La bibliographie de la partie ”Transformée de Fourier irrégulière”

A. Coulais

9 avril 2003

Références

Bart95 [1] D. Barthès, Cleaning power spectra of short, unevenly spaced time-series. application to LPVs, A&A Supplement

Series, 111(0) :373–386, Juin 1995, .

COB92 [2] M. Carbonell, R. Oliver, and J. Ballester, Power spectra of gapped time series : a comparison of several

methods, A&A, 264(0) :350–360, mois inconnu 1992, .

CPL85 [3] J. Clark, M. Palmer, and P. Lawrence, A transformation method for the reconstruction of functions from nonuniform spaced samples, IEEE Trans. on Acoustics, Speech, and Signal Processing, 33(4) :1151–1165, Octobre 1985, .

Fost95 [4] G. Foster, The cleanest Fourier spectrum, The Astronomical Journal, 109(4) :1889, Avril 1995, . Fost96 [5] G. Foster, Wavelets for period analysis of unevenly sampled time series, The Astronomical Journal, 112(4) :1709–

1729, Octobre 1996, .

Jer77 [6] A. Jerri, The Shannon sampling theorem. Its various extensions and applications : a tutorial review, Proc. IEEE,

65(11) :1565–1596, Novembre 1977, .

KiBo90 [7] S. Kim and N. Bose, Reconstruction of 2-D bandlimited discrete signals from nonuniform samples, Proc. IEEE,

137, Pt. F(3) :197–204, Juin 1990, .

Kuhn82 [8] J. Kuhn, Recovering spectral information from unevenly sampled data : two machine-efficient solutions, The Astro-

nomical Journal, 87(1) :196–202, Janvier 1982, .

Leg93 [9] C. Léger, R. Lédée, J. Thiel, L. Pourcelot, and F. Patat, Restitution complète d’un contour non fermé du

ventricule gauche par le modèle SFDP, journal inconnu, 0(0) :0–0, mois inconnu 1993, .

LuMa92 [10] P. Luchini and F. Manzo, A fast conformal mapping algorithm with no FFT, Journal of Computational Physics,

101(0) :368–374, mois inconnu 1992, .

Papou66 [11] A. Papoulis, Error analysis in sampling theory, Proc. IEEE, 54 :947–955, Juillet 1966, . Shaw71 [12] L. Shaw, Spectral estimates from nonuniform samples, IEEE Trans. on Audio and Electroacoustics, 19(1) :24–31,

Mars 1971, .

Sou88 [13] M. Soumekh, Band-limited interpolation from unevenly spaced sampled data, IEEE Trans. on Acoustics, Speech,

and Signal Processing, 36(1) :110–122, Janvier 1988, .

Swan82 [14] P. Swan, Discrete Fourier transforms of nonuniformly spaced data, The Astronomical Journal, 87(11) :1608–1615,

Novembre 1982, .

Yen56 [15] J. Yen, On the nonuniform sampling of bandwidth-limited signals, IRE Trans. Circuit Theory, CT–3 :251–257, Décembre 1956, .